# PairFlow: Closed-Form Source-Target Coupling for Few-Step Generation in Discrete Flow Models

### v1

@@ 1 Introduction
Discrete Flow Models (DFMs) (Campbell:2024MultiFlow; Gat:2024DFM) have recently emerged as a promising class of generative models, extending the idea of Flow Models (FMs)  for continuous data to the discrete domain. By adapting flow-based principles to categorical structures, DFMs provide a principled and efficient way to capture complex discrete distributions through iterative sampling. They have shown success across a variety of applications, particularly in scientific domains such as molecule generation (ramakrishnan2014quantum; irwin2012zinc), where DFMs offer a natural framework for modeling chemical structures and generating novel candidates.
Analogous to FMs in the continuous domain, a key challenge of DFMs is the long computation time for generation due to their iterative sampling nature. Recent work (Deschenaux:2025sdtt; Hayakawa:2025di4c; Sahoo:2025Duo; Yoo:2025ReDi) have sought to accelerate the generative process through distillation-based finetuning, which builds on ideas originally developed for continuous flow matching. Notably, ReFlow (Liu:2023RF) is a well-known technique for FMs that pairs samples from the source (prior) distribution and the target distribution by running the generative process of a pretrained model and using the resulting pairs for finetuning. Recently, this idea has also been extended to DFMs (Yoo:2025ReDi) to reduce conditional total correlation through finetuning, thereby enabling few-step generation.
@@
At the core of our framework is the simulation of probability paths connecting source (prior) and target (data) distributions in discrete spaces, made possible by closed-form expressions of velocities. While closed-form forward velocities have been studied for flow models in continuous domains (Karras:2022EDM; Bertrand:2025Closed), they have, to the best of our knowledge, neither been explored for DFMs nor applied to identifying suitable source–target pairs in the context of distillation-based acceleration, as in ReDi (Yoo:2025ReDi). In this work, we investigate this idea for the first time. For DFMs, with a particular focus on uniform-state models (Sahoo:2025Duo; Schiff:2025UDLM) equipped with a self-correcting mechanism, we show that the closed-form forward velocity is determined by the Hamming distance, which measures the number of differing tokens between two sequences. Using this velocity, samples from the source (latent) distribution can be mapped to given target samples. However, because multiple source samples may map to the same target, covering all targets through coupling would require an impractically large number of source samples. To overcome this, we derive the corresponding backward velocity in closed form and leverage it to simulate backward probability paths that efficiently map data points to source points, making pair discovery computationally efficient.
In our experiments, we show that the proposed framework enables few-step sampling across diverse discrete domains, including molecular data (ramakrishnan2014quantum; irwin2012zinc) and 2D images, exemplified by MNIST-Binary (lecun2002gradient) and CIFAR-10 (krizhevsky2009learning). On the QM9 (ramakrishnan2014quantum) and ZINC-250k (irwin2012zinc) datasets,  PairFlow not only improves the base model but also performs comparably to, or even better than, distilled models that require up to $143\times$ more compute during finetuning, compared to our lightweight preprocessing algorithm. Similar improvements are observed on MNIST-Binary, where models paired with PairFlow achieve performance comparable to those using DCD (Sahoo:2025Duo) and ReDi (Yoo:2025ReDi), while being up to $35\times$ faster. Furthermore, after subsequent distillation, base models trained with pairs generated by our method consistently achieve higher performance, underscoring the importance of well-constructed source–target pairings.
@@ 2.1 Discrete Flow models
The concept of Flow Matching (Lipman:2023FlowMatching) has recently been extended to the discrete flow-based models (Gat:2024DFM; Campbell:2024MultiFlow; Sahoo:2024MDLM; Schiff:2025UDLM), demonstrating its flexibility across high-dimensional and structured data (Bai:2025meissonic; Chang:2022maskgit; Weber:2024maskbit; Arriola:2025blockdiffusion; Nie:2025largeldm; Yu:2023magvit; Lee:2025genmol; Campbell:2024MultiFlow; Wang:2025drakes). Among these, uniform-state models (Sahoo:2025Duo; Schiff:2025UDLM) have been studied for their self-correcting properties, which enable recovery from errors introduced during parallel decoding. However, their performance degrades markedly in few-step settings, posing a key limitation for efficient generation under tight compute budgets.
@@ 2.2 Accelerating Discrete Flow models
Several approaches have been proposed to accelerate sampling with DFMs. Park:2025jumpyoursteps directly optimize sampling timesteps for improved parallelism while mitigating decoding errors. Hayakawa:2025di4c highlight the importance of capturing dimensional correlations for faster sampling and introduce mixture models to this end, at the cost of additional loss terms that complicate optimization. Another line of work adapts techniques from continuous domains, as in Sahoo:2025Duo, that propose a discrete analogue of Consistency Distillation (CD) (Song:2023consistency) by leveraging the duality between uniform-state and continuous Gaussian models. Most relevant to our approach is ReDi (Yoo:2025ReDi), which draws inspiration from the concept of straight flows in ReFlow (Liu:2023RF) and iteratively optimizes pairs of data and noise samples.
@@ 3 Preliminaries
In this section, we provide a brief overview of flow matching for generative modeling of discrete data (Sec. 3.1), followed by a rectification technique (Yoo:2025ReDi) that enables faster generation (Sec. 3.2) by reducing total correlation errors.
@@ 3.1 Discrete Flow matching
The goal of Discrete Flow Matching (DFM) (Campbell:2024MultiFlow; Gat:2024DFM) is to learn a probability path $p_{t}(\cdot)$ that connects a known, easy-to-sample source distribution $p(x)$ to an unknown target distribution $q(x)$, both defined over a discrete state space. Once $p_{t}(\cdot)$ is known, samples from $q$ can be generated by drawing $x_{0}\sim p$ and transporting it along the path.
@@
Given a target probability path $p_{t}(x):\mathcal{V}^{N}\times[0,1]\to[0,1]$ with an associated velocity field $v_{t}(x):\mathcal{V}^{N}\times[0,1]\to\mathbb{R}^{N\times K}$, we introduce a network $p_{1|t}^{\theta}(x):\mathcal{V}^{N}\times[0,1]\to(\Delta^{K})^{N}$ to approximate $v_{t}(x)$. Its parameters $\theta$ are optimized via the DFM objective (Gat:2024DFM):
@@
As token-wise conditional paths $p_{t}(z^{i}|x_{0},x_{1})$Gat:2024DFM employ the mixture path of form:
@@ 3.2 Straightening Probability Paths for Accelerated Sampling
The concept of straight probability paths was originally introduced in the continuous domain to enable accelerated sampling. Prior work (liuflow) identified curved probability paths as a key challenge in few-step sampling: when velocity fields are evaluated only at coarse time steps, numerical integration deviates from the true trajectories. liuflow addressed this issue through “rectification,” in which a student flow model is trained on source–target pairs generated by a teacher model, effectively yielding significantly straighter probability paths.
In the discrete setting, this challenge of path curvature translates to capturing the statistical correlations between tokens. Since DFMs approximate exponentially large joint transitions through fully factorized per-token updates, a mismatch inevitably arises between the true joint transition and its product-form approximation. This discrepancy becomes especially detrimental during few-step generation, where highly correlated tokens must be updated simultaneously. To address this, prior works have primarily relied on distillation-based approaches (Hayakawa:2025di4c; Sahoo:2025Duo; Deschenaux:2025sdtt), aiming to better capture these correlations by explicitly transferring multi-step dependencies from a stronger teacher model.
Yoo:2025ReDi formalized this factorization mismatch via conditional Total Correlation (TC), defined as:
@@
which serves as a metric for the factorization error. Crucially, Yoo:2025ReDi interpret this factorization error as the discrete analog of path curvature: minimizing TC is equivalent to “straightening” the trajectory by decoupling token transitions. Analogous to ReFlow (liuflow), which rectifies continuous paths, they demonstrate that reducing Eqn. 7 requires iteratively refining the source–target coupling $\pi(x_{0},x_{1})$. To achieve this, they employ an iterative distillation process, alternating between generating improved pairs using the current model and optimizing $\mathcal{L}_{\text{DFM}}$. This procedure effectively finds a “statistically straight” coupling that enables efficient few-step generation.
@@ 4 PairFlow
For DFMs, ReDi (Yoo:2025ReDi) improves sample quality in few-step generation by rectifying source-target pairs. However, it relies on samples from a pretrained model followed by costly retraining or finetuning. We take this one step further and pose the following question: What if these pairs could be generated directly from the data, without relying on a pretrained model or sampling from it?
To address this question, we propose a principled approach for discovering well-aligned source–target pairs without relying on pretrained models, enabling models trained on such pairs to achieve strong performance with few-step sampling. Our method, termed PairFlow, leverages closed-form velocity fields that can be computed directly from the data samples, requiring only prior knowledge of the source distribution. We assume this distribution to be uniform, a choice extensively studied in recent work (Sahoo:2025Duo; Schiff:2025UDLM), as models trained under this prior naturally acquire self-correcting properties.
In Sec. 4.1, we introduce the closed-form forward velocity for discrete flow matching (Gat:2024DFM). In Sec. 4.2, we extend this to the closed-form backward velocity and propose an algorithm for discovering well-aligned source–target pairs during the preprocessing phase.
@@ 4.1 Finding Pairs via Closed-Form Forward Velocity Fields
As discussed in Sec. 3.1, DFMs (Campbell:2024MultiFlow; Gat:2024DFM) aim to learn a marginal velocity field $v_{t}(\cdot)$ that induces a probability path $p_{t}(\cdot)$, transporting the source distribution $p_{0}(\cdot)$ to the target distribution $q(\cdot)=p_{1}(\cdot)$, which is unknown in practice. Instead, we only have access to a finite dataset of $M$ samples $\{d_{m}\}_{m=1}^{M}$. This empirical distribution $\tilde{q}(x)$ can be represented as a mixture of Dirac deltas centered at the observed samples:
@@
For continuous domains, Karras:2022EDM; Bertrand:2025Closed have shown that the velocity field transporting $p_{0}$ to $q$ can be derived in closed form when both distributions admit tractable density expressions. To the best of our knowledge, this idea has not been explored in discrete domains; in the following, we derive the closed-form velocity field for discrete domains for the first time.
@@ 4.2 Finding Pairs via Closed-Form Backward Velocity Fields
We address this issue by backtracing trajectories along $p_{t}(\cdot)$, starting from $\tilde{q}(x)$ and progressing toward the source distribution $p_{0}(\cdot)$. Unlike the forward construction in Sec. 4.1, this guarantees that all data points in $\tilde{q}(x)$ are included in the resulting pairs by design. As illustrated at the top of Fig. 1PairFlow inverts data samples toward the source distribution, assumed to be uniform. Unlike the standard corruption process used by UDLM (Schiff:2025UDLM) shown at the bottom of Fig. 1, the source samples obtained by PairFlow remain closer to the original data in terms of Hamming distance. This helps the model learn to recover data with fewer token transitions during training, effectively approximating the straight probability paths explored in ReFlow (Liu:2023RF) and ReDi (Yoo:2025ReDi).
@@
with a detailed derivation provided in App. A.2. Substituting this into the definition of the backward velocity field from Gat:2024DFM
@@ 5.1 Experiment Setup
Across multiple benchmarks, we compare our approach against state-of-the-art discrete flow models, including MDLM (Sahoo:2024MDLM) and UDLM (Schiff:2025UDLM). Since our method is based on a uniform source distribution, we adopt UDLM (Schiff:2025UDLM), the leading uniform-state model, as our base and denote UDLM trained with pairs generated by Alg. 1 as PairFlow throughout the remainder of this section. In addition, we compare against these models augmented with distillation-based techniques that require additional finetuning, Discrete Consistency Distillation (DCD) (Sahoo:2025Duo) and ReDi (Yoo:2025ReDi), denoted throughout this section by the suffixes “+ DCD” and “+ ReDi”. The detailed training setups of these models, such as hyperparameters, are provided in App. B. Additionally, we report the performance of the same base model trained on pairs formed by each data point and a source sample randomly drawn from the uniform distribution with our detailed experimental results in App. D.
@@
We evaluate our method across a diverse set of discrete generative modeling benchmarks, covering both molecule and image generation tasks. For molecule generation, we experiment with the QM9 (ramakrishnan2014quantum) and ZINC-250k (irwin2012zinc) datasets. For image generation, we use the MNIST dataset (lecun2002gradient) with binarized pixel values (denoted MNIST-Binary) and the CIFAR-10 dataset (krizhevsky2009learning), where pixel intensities are scaled to 8-bit integers, and horizontal flip augmentation is applied. Dataset statistics, including sample size, vocabulary size, and overall dataset size, are summarized in Tab. 1.
@@
For molecular generation, we follow Schiff:2025UDLM and evaluate the validity, uniqueness, and novelty of generated molecules. Specifically, we sample 1,024 SMILES strings (weininger1988smiles), convert them into molecular graphs, and compute these metrics. All results are averaged over 10 trials, with further details provided in App. D. For image generation, we report Fréchet Inception Distance (FID) (Heusel:2017FID) and Inception Score (IS) (Salimans:2016IS). FID is computed with 1,000 images for MNIST-Binary, and both FID and IS are computed with 5,000 generated images for CIFAR-10. The training dataset is used as the reference for FID computation. Across all experiments, we vary the number of sampling steps to evaluate performance in both low- and high-NFE settings. In particular, we generate samples using 1641-64 steps for molecular benchmarks (QM9 and ZINC-250k) and MNIST-Binary benchmark, and 810248-1024 steps for CIFAR-10 (krizhevsky2009learning), as models yielded excessively high FIDs under extremely low-step settings.

### v2

@@ 1 Introduction
Discrete Flow Models (DFMs) (Campbell et al., 2024; Gat et al., 2024) have recently emerged as a promising class of generative models, extending the idea of Flow Models (FMs)  for continuous data to the discrete domain. By adapting flow-based principles to categorical structures, DFMs provide a principled and efficient way to capture complex discrete distributions through iterative sampling. They have shown success across a variety of applications, particularly in scientific domains such as molecule generation (Ramakrishnan et al., 2014; Irwin et al., 2012), where DFMs offer a natural framework for modeling chemical structures and generating novel candidates.
Analogous to FMs in the continuous domain, a key challenge of DFMs is the long computation time for generation due to their iterative sampling nature. Recent work (Deschenaux and Gulcehre, 2025; Hayakawa et al., 2025; Sahoo et al., 2025; Yoo et al., 2025) have sought to accelerate the generative process through distillation-based finetuning, which builds on ideas originally developed for continuous flow matching. Notably, ReFlow (Liu et al., 2023) is a well-known technique for FMs that pairs samples from the source (prior) distribution and the target distribution by running the generative process of a pretrained model and using the resulting pairs for finetuning. Recently, this idea has also been extended to DFMs (Yoo et al., 2025) to reduce conditional total correlation through finetuning, thereby enabling few-step generation.
@@
At the core of our framework is the simulation of probability paths connecting source (prior) and target (data) distributions in discrete spaces, made possible by closed-form expressions of velocities. While closed-form forward velocities have been studied for flow models in continuous domains (Karras et al., 2022; Bertrand et al., 2025), they have, to the best of our knowledge, neither been explored for DFMs nor applied to identifying suitable source–target pairs in the context of distillation-based acceleration, as in ReDi (Yoo et al., 2025). In this work, we investigate this idea for the first time. For DFMs, with a particular focus on uniform-state models (Sahoo et al., 2025; Schiff et al., 2025) equipped with a self-correcting mechanism, we show that the closed-form forward velocity is determined by the Hamming distance, which measures the number of differing tokens between two sequences. Using this velocity, samples from the source (latent) distribution can be mapped to given target samples. However, because multiple source samples may map to the same target, covering all targets through coupling would require an impractically large number of source samples. To overcome this, we derive the corresponding backward velocity in closed form and leverage it to simulate backward probability paths that efficiently map data points to source points, making pair discovery computationally efficient.
In our experiments, we show that the proposed framework enables few-step sampling across diverse discrete domains, including molecular data (Ramakrishnan et al., 2014; Irwin et al., 2012) and 2D images, exemplified by MNIST-Binary (LeCun et al., 2002) and CIFAR-10 (Krizhevsky and others, 2009). On the QM9 (Ramakrishnan et al., 2014) and ZINC-250k (Irwin et al., 2012) datasets,  PairFlow not only improves the base model but also performs comparably to, or even better than, distilled models that require up to $143\times$ more compute during finetuning, compared to our lightweight preprocessing algorithm. Similar improvements are observed on MNIST-Binary, where models paired with PairFlow achieve performance comparable to those using DCD (Sahoo et al., 2025) and ReDi (Yoo et al., 2025), while being up to $35\times$ faster. Furthermore, after subsequent distillation, base models trained with pairs generated by our method consistently achieve higher performance, underscoring the importance of well-constructed source–target pairings.
@@ 2.1 Discrete Flow models
The concept of Flow Matching (Lipman et al., 2023) has recently been extended to the discrete flow-based models (Gat et al., 2024; Campbell et al., 2024; Sahoo et al., 2024; Schiff et al., 2025), demonstrating its flexibility across high-dimensional and structured data (Bai et al., 2025; Chang et al., 2022; Weber et al., 2024; Arriola et al., 2025; Nie et al., ; Yu et al., 2023; Lee et al., 2025; Campbell et al., 2024; Wang et al., 2025). Among these, uniform-state models (Sahoo et al., 2025; Schiff et al., 2025) have been studied for their self-correcting properties, which enable recovery from errors introduced during parallel decoding. However, their performance degrades markedly in few-step settings, posing a key limitation for efficient generation under tight compute budgets.
@@ 2.2 Accelerating Discrete Flow models
Several approaches have been proposed to accelerate sampling with DFMs. Park et al. (2024) directly optimize sampling timesteps for improved parallelism while mitigating decoding errors. Hayakawa et al. (2025) highlight the importance of capturing dimensional correlations for faster sampling and introduce mixture models to this end, at the cost of additional loss terms that complicate optimization. Another line of work adapts techniques from continuous domains, as in Sahoo et al. (2025), that propose a discrete analogue of Consistency Distillation (CD) (Song et al., 2023) by leveraging the duality between uniform-state and continuous Gaussian models. Most relevant to our approach is ReDi (Yoo et al., 2025), which draws inspiration from the concept of straight flows in ReFlow (Liu et al., 2023) and iteratively optimizes pairs of data and noise samples.
@@ 3 Preliminaries
In this section, we provide a brief overview of flow matching for generative modeling of discrete data (Sec. 3.1), followed by a rectification technique (Yoo et al., 2025) that enables faster generation (Sec. 3.2) by reducing total correlation errors.
@@ 3.1 Discrete Flow matching
The goal of Discrete Flow Matching (DFM) (Campbell et al., 2024; Gat et al., 2024) is to learn a probability path $p_{t}(\cdot)$ that connects a known, easy-to-sample source distribution $p(x)$ to an unknown target distribution $q(x)$, both defined over a discrete state space. Once $p_{t}(\cdot)$ is known, samples from $q$ can be generated by drawing $x_{0}\sim p$ and transporting it along the path.
@@
Given a target probability path $p_{t}(x):\mathcal{V}^{N}\times[0,1]\to[0,1]$ with an associated velocity field $v_{t}(x):\mathcal{V}^{N}\times[0,1]\to\mathbb{R}^{N\times K}$, we introduce a network $p_{1|t}^{\theta}(x):\mathcal{V}^{N}\times[0,1]\to(\Delta^{K})^{N}$ to approximate $v_{t}(x)$. Its parameters $\theta$ are optimized via the DFM objective (Gat et al., 2024):
@@
As token-wise conditional paths $p_{t}(z^{i}|x_{0},x_{1})$Gat et al. (2024) employ the mixture path of form:
@@ 3.2 Straightening Probability Paths for Accelerated Sampling
The concept of straight probability paths was originally introduced in the continuous domain to enable accelerated sampling. Prior work (Liu et al., 2023) identified curved probability paths as a key challenge in few-step sampling: when velocity fields are evaluated only at coarse time steps, numerical integration deviates from the true trajectories. Liu et al. (2023) addressed this issue through “rectification,” in which a student flow model is trained on source–target pairs generated by a teacher model, effectively yielding significantly straighter probability paths.
In the discrete setting, this challenge of path curvature translates to capturing the statistical correlations between tokens. Since DFMs approximate exponentially large joint transitions through fully factorized per-token updates, a mismatch inevitably arises between the true joint transition and its product-form approximation. This discrepancy becomes especially detrimental during few-step generation, where highly correlated tokens must be updated simultaneously. To address this, prior works have primarily relied on distillation-based approaches (Hayakawa et al., 2025; Sahoo et al., 2025; Deschenaux and Gulcehre, 2025), aiming to better capture these correlations by explicitly transferring multi-step dependencies from a stronger teacher model.
Yoo et al. (2025) formalized this factorization mismatch via conditional Total Correlation (TC), defined as:
@@
which serves as a metric for the factorization error. Crucially, Yoo et al. (2025) interpret this factorization error as the discrete analog of path curvature: minimizing TC is equivalent to “straightening” the trajectory by decoupling token transitions. Analogous to ReFlow (Liu et al., 2023), which rectifies continuous paths, they demonstrate that reducing Eqn. 7 requires iteratively refining the source–target coupling $\pi(x_{0},x_{1})$. To achieve this, they employ an iterative distillation process, alternating between generating improved pairs using the current model and optimizing $\mathcal{L}_{\text{DFM}}$. This procedure effectively finds a “statistically straight” coupling that enables efficient few-step generation.
@@ 4 PairFlow
For DFMs, ReDi (Yoo et al., 2025) improves sample quality in few-step generation by rectifying source-target pairs. However, it relies on samples from a pretrained model followed by costly retraining or finetuning. We take this one step further and pose the following question: What if these pairs could be generated directly from the data, without relying on a pretrained model or sampling from it?
To address this question, we propose a principled approach for discovering well-aligned source–target pairs without relying on pretrained models, enabling models trained on such pairs to achieve strong performance with few-step sampling. Our method, termed PairFlow, leverages closed-form velocity fields that can be computed directly from the data samples, requiring only prior knowledge of the source distribution. We assume this distribution to be uniform, a choice extensively studied in recent work (Sahoo et al., 2025; Schiff et al., 2025), as models trained under this prior naturally acquire self-correcting properties.
In Sec. 4.1, we introduce the closed-form forward velocity for discrete flow matching (Gat et al., 2024). In Sec. 4.2, we extend this to the closed-form backward velocity and propose an algorithm for discovering well-aligned source–target pairs during the preprocessing phase.
@@ 4.1 Finding Pairs via Closed-Form Forward Velocity Fields
As discussed in Sec. 3.1, DFMs (Campbell et al., 2024; Gat et al., 2024) aim to learn a marginal velocity field $v_{t}(\cdot)$ that induces a probability path $p_{t}(\cdot)$, transporting the source distribution $p_{0}(\cdot)$ to the target distribution $q(\cdot)=p_{1}(\cdot)$, which is unknown in practice. Instead, we only have access to a finite dataset of $M$ samples $\{d_{m}\}_{m=1}^{M}$. This empirical distribution $\tilde{q}(x)$ can be represented as a mixture of Dirac deltas centered at the observed samples:
@@
For continuous domains, Karras et al. (2022); Bertrand et al. (2025) have shown that the velocity field transporting $p_{0}$ to $q$ can be derived in closed form when both distributions admit tractable density expressions. To the best of our knowledge, this idea has not been explored in discrete domains; in the following, we derive the closed-form velocity field for discrete domains for the first time.
@@ 4.2 Finding Pairs via Closed-Form Backward Velocity Fields
We address this issue by backtracing trajectories along $p_{t}(\cdot)$, starting from $\tilde{q}(x)$ and progressing toward the source distribution $p_{0}(\cdot)$. Unlike the forward construction in Sec. 4.1, this guarantees that all data points in $\tilde{q}(x)$ are included in the resulting pairs by design. As illustrated at the top of Fig. 1PairFlow inverts data samples toward the source distribution, assumed to be uniform. Unlike the standard corruption process used by UDLM (Schiff et al., 2025) shown at the bottom of Fig. 1, the source samples obtained by PairFlow remain closer to the original data in terms of Hamming distance. This helps the model learn to recover data with fewer token transitions during training, effectively approximating the straight probability paths explored in ReFlow (Liu et al., 2023) and ReDi (Yoo et al., 2025).
@@
with a detailed derivation provided in App. A.2. Substituting this into the definition of the backward velocity field from Gat et al. (2024)
@@ 5.1 Experiment Setup
Across multiple benchmarks, we compare our approach against state-of-the-art discrete flow models, including MDLM (Sahoo et al., 2024) and UDLM (Schiff et al., 2025). Since our method is based on a uniform source distribution, we adopt UDLM (Schiff et al., 2025), the leading uniform-state model, as our base and denote UDLM trained with pairs generated by Alg. 1 as PairFlow throughout the remainder of this section. In addition, we compare against these models augmented with distillation-based techniques that require additional finetuning, Discrete Consistency Distillation (DCD) (Sahoo et al., 2025) and ReDi (Yoo et al., 2025), denoted throughout this section by the suffixes “+ DCD” and “+ ReDi”. The detailed training setups of these models, such as hyperparameters, are provided in App. B. Additionally, we report the performance of the same base model trained on pairs formed by each data point and a source sample randomly drawn from the uniform distribution with our detailed experimental results in App. D.
@@
We evaluate our method across a diverse set of discrete generative modeling benchmarks, covering both molecule and image generation tasks. For molecule generation, we experiment with the QM9 (Ramakrishnan et al., 2014) and ZINC-250k (Irwin et al., 2012) datasets. For image generation, we use the MNIST dataset (LeCun et al., 2002) with binarized pixel values (denoted MNIST-Binary) and the CIFAR-10 dataset (Krizhevsky and others, 2009), where pixel intensities are scaled to 8-bit integers, and horizontal flip augmentation is applied. Dataset statistics, including sample size, vocabulary size, and overall dataset size, are summarized in Tab. 1.
@@
For molecular generation, we follow Schiff et al. (2025) and evaluate the validity, uniqueness, and novelty of generated molecules. Specifically, we sample 1,024 SMILES strings (Weininger, 1988), convert them into molecular graphs, and compute these metrics. All results are averaged over 10 trials, with further details provided in App. D. For image generation, we report Fréchet Inception Distance (FID) (Heusel et al., 2017) and Inception Score (IS) (Salimans et al., 2016). FID is computed with 1,000 images for MNIST-Binary, and both FID and IS are computed with 5,000 generated images for CIFAR-10. The training dataset is used as the reference for FID computation. Across all experiments, we vary the number of sampling steps to evaluate performance in both low- and high-NFE settings. In particular, we generate samples using 1641-64 steps for molecular benchmarks (QM9 and ZINC-250k) and MNIST-Binary benchmark, and 810248-1024 steps for CIFAR-10 (Krizhevsky and others, 2009), as models yielded excessively high FIDs under extremely low-step settings.